Shift report 20010824 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Aug 24 2001 - 16:08:05 EDT

  • Next message: Apache: "Shift report 20010824 16:00-24:00"

    ------------------------------------------------------------
    Date: Fri Aug 24 16:08:05 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#8 pages 15 - 17
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam is stored and cogged around 14:45.
       Now 56 bunches.
    
    2. Run 4361 : Pedestal Run.
       Run 4362 : Data Run, 126K trigger 1.
                  T5 was not working. (see supermon!)
       Run 4363 : Now running.     
    
    3. BRAT-2-0-32, and BROP-1-0-5 installed this morning.
       Dana changed ZDC voltages, not threshold.
       Dana wants to see "neutron peak!".
       T5 Low voltages recycled. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Aug 24 2001 - 16:08:49 EDT